The ages of two friends is in the ratio 5:6. After how many years will the ages be in the ratio 7:8?
i want solution is steps .......please

Answer by Jk22(389)

Suppose the ratio's numbers are the ages directly :
we notice that 7=5+2 and 8=6+2, so if b was 6 years, it will happen after 2 years.
Suppose the ratios were already simplified, and hence multiply by a fixed number a the first ratio : 5a/6a and the second : 7a/8a, the number of years elapsed is :
7a-5a=2a for the first, 8a-6a=2a for the second.
Hence the time elapsed were in this case 2a, with a any non zero number, hence it can be any time !